Which of the following points lies on the line 5y-2x=-10?
(0,5)
(5,0)
(2,0) or
(0,2)
Thank you!!

Answer by Alan3354(30993)

Which of the following points lies on the line 5y-2x=-10?
(0,5)
(5,0)
(2,0) or
(0,2)
---------
5y-2x=-10
----------------
2x - 5y = 10
Try (0,5)
Sub 0 for x and 5 for y
2*0 - 5*5 =? 10
Doesn't work.
---------
Try the others the same way.